OpenStudy (anonymous):

f(x)=-3-x, g(y)= square root of y+9 g(f(x))= square root of -3-x+9 what is the domain of the function of g(f(x))?

OpenStudy (campbell_st):

\[g(f(x) = \sqrt{-3-x + 9}...or .....g(f(x) = \sqrt{6 - x}\] so radicals can't be negative, so for the domain you need to solve \[6 - x \ge 0\] and you can always check by picking a value in your domain and testing.

OpenStudy (anonymous):

this how my problem looks like \[f(x)=-3-x, g(y)=\sqrt{y+9} g(f(x))=\sqrt{-3-x+9}\]

OpenStudy (anonymous):

Guys. Before you finish anything, the equation is: \[g(y) = \sqrt{y} + 9\] I know you said that +9 is included, but then your options would be invalid.

OpenStudy (mathstudent55):

Thanks, now I can unsderstand it. Question: \(f(x)=-3-x\) \(g(y)= \sqrt{y+9} \) \( g(f(x))= \sqrt{-3-x+9 } \) What is the domain of the function g(f(x))? Solution: g(f(x)) can be simplified to: \( g(f(x)) = \sqrt{6 - x} \) Since there is a root in the comopsition, we need to avoid negative values of the radicand. \(6 - x \ge 0 \) \(-x \ge -6\) \(x \le 6\) This tells you that the domain of the compositiin is all numbers less than or equal to 6.

OpenStudy (anonymous):

yeah ur right damn i type too fast that i sometimes make mistakes my bad

OpenStudy (anonymous):

a. all real numbers less than or equal to 3 b. all numbers greater than or equal to 3 c. all numbers less than or equal to -3 d. all numbers greater than or equal to -3

OpenStudy (anonymous):

yes this are my options

OpenStudy (anonymous):

i think its a

OpenStudy (mathstudent55):

None of your choices have the correct answer. Are you sure these are the correct choices for this problem?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

a. all numbers less than or equal to 3 b. all numbers greater than or equal to 3 c. all numbers less than or equal to -3 d. all numbers greater than or equal to -3

OpenStudy (anonymous):

DOMAIN: x ≤ -3 http://www.wolframalpha.com/input/?i= √%28-3-x%29+%2B+9

OpenStudy (anonymous):

so c

OpenStudy (mathstudent55):

@romanortiz65 Is \( g(y) = \sqrt{y} + 9 \) or \(g(y) = \sqrt{y + 9} \)

OpenStudy (anonymous):

its the first choice u wrote

OpenStudy (anonymous):

\[g(y)=\sqrt{y}+9\]

OpenStudy (anonymous):

hello anyone there that can still help me

OpenStudy (anonymous):

so the answer is C

OpenStudy (anonymous):

so its c

OpenStudy (anonymous):

so its c

OpenStudy (mathstudent55):

That's not what you wrote above. No wonder my solution didn't match any of the choices.

OpenStudy (mathstudent55):

In your first response after you opened the post, you show (y + 9) inside the square root. Now you are saying only y should be inside the square root.

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

a. all numbers less than or equal to 3 b. all numbers greater than or equal to 3 c. all numbers less than or equal to -3 d. all numbers greater than or equal to -3

OpenStudy (anonymous):

Yes. Sorry! I wasn't at my computer...